# Heights and periods of algebraic cycles in families

### v1

@@ 3. The Betti rank and Zariski closedness of the rank-strata
The Betti rank is given by the following formula:
@@
For each $t\geq 0$, the set $S_{{\mathscr{F}}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$. In particular, $r(\nu)=\dim S-\min\{t\geq 0:S_{{\mathscr{F}}}(t)=S\}$.
The proofs of Theorem 3.1 and Theorem 3.2 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the case when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable. A key ingredient for our proof is the mixed Ax-Schanuel theorem, which is used multiple times. In [Gao20a], the version for universal abelian varieties [Gao20b, Thm. 1.1] was used. In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, for two reasons. First, it eases the notation for the proof. Second and more importantly, [BBT23] is necessary to prove the full version of Theorem 1.6. Still, we point out that [BBT23] is not needed in the case which we care the most about in this paper, i.e. $r(\nu)=\dim S$ if and only if the RHS of (1.6) equals $\dim S$ because we can easily reduce the proof to the case where every fiber of $\varphi$ has dimension 00.
@@ 3.1. Replacing $S$ by $\varphi(S)$
(3.3) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
@@ 3.2. Bi-algebraic system and Ax–Schanuel
    Let $\tilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset. The weakly special closure of $\tilde{Y}$, denoted by $\tilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\tilde{Y}$.
@@
The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi21] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
@@
Let $\tilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
(3.4) $\dim\tilde{Z}^{\mathrm{Zar}}+\dim u(\tilde{Z})^{\mathrm{Zar}}\geq\dim\tilde{Z}^{\mathrm{ws}}+\dim\tilde{Z},$
where $\tilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\tilde{Z}$.
@@
Let $Y:=u(\tilde{Z})^{\mathrm{Zar}}$. Let $\mathscr{Z}:=\{(z,y)\in\tilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$. The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\tilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\tilde{Z}$. Then (3.4) is a direct consequence of the mixed Ax–Schanuel theorem [GK24, Thm. 1.1] applied to $\mathscr{Z}$. ∎
@@
An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds true: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
@@ 3.3. Applications of Ax–Schanuel
For each $t\geq 0$, $S_{{\mathscr{F}}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
@@
    $S_{{\mathscr{F}}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
@@
Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S_{{\mathscr{F}}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
$Y:=\overline{u(\{a\}\times\tilde{C})}^{\mathrm{Zar}},$
for some complex analytic irreducible $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}=t$ and some $a\in V({\mathbb{R}})$.
Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\tilde{C}$). Then we get
$\dim\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}+t.$
By Lemma B.4, $\overline{\{a\}\times\tilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\tilde{C}}^{\mathrm{Zar}}$. Hence
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\overline{\tilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
The last inequality holds because $\overline{\tilde{C}}^{\mathrm{Zar}}\subseteq\tilde{C}^{\mathrm{ws}}$. So
$\dim Y\geq\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}-\dim\tilde{C}^{\mathrm{ws}}+t.$
Now, to prove (i), it suffices to prove $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
Let us prove $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\tilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$. By definition of $Y$, we have $u(\{a\}\times\tilde{C})\subseteq Y$. Hence $u((\{a\}\times\tilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$. On the other hand, $u((\{a\}\times\tilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\tilde{C})^{\mathrm{ws}})$. Now we have established $u((\{a\}\times\tilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
Similarly, we have $\dim\tilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
@@
There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds true. For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}$ is the image of an $N_{j}({\mathbb{R}})^{+}$-orbit contained in ${\mathscr{D}}^{\prime}_{j}$ under $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ for some $j\in\{1,\ldots,k\}$.
The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds true without replacing $S$ by $\varphi(S)$.
@@
$S_{{\mathscr{F}}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
@@
Let $t\geq 0$. By Proposition 3.7, $S_{{\mathscr{F}}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$. Then by Proposition 3.8, $Y^{\mathrm{ws}}$ is a fiber $[p_{N_{j}}]$ for some $j\in\{1,\ldots,k\}$, and hence $\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})=\frac{1}{2}\dim(V\cap N_{j})$. So $\dim Y\geq\frac{1}{2}\dim(V\cap N_{j})+t$. So $Y\subseteq E_{j}(t)$ because $[p_{N_{j}}](Y)$ is a point. ∎
@@
Let us prove $\leq$”. By (2.7) we have $r(\nu)=\max_{\tilde{s}\in\tilde{S}}(\mathrm{d}\tilde{\varphi}_{V})_{\tilde{s}}$. Hence $r(\nu)\leq\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)$ for any normal subgroup $N\lhd\mathbf{G}$.
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S_{{\mathscr{F}}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
$S_{{\mathscr{F}}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
@@
If $S=S_{\geq t}$, then $\geq$” holds true already for $N=\{1\}$.
@@ 3.5. Zariski closedness of the degeneracy loci
For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S_{{\mathscr{F}}}(t)$.
@@
For each $s\in E_{j}(t)$, by definition there exist an irreducible $\tilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\tilde{Y})$, $\dim\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\tilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$. The last condition implies that $\tilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$. Hence by the discussion above, $\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$. Now that $\dim_{{\mathbb{C}}}\tilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds true: For each $(a,x_{0})\in\tilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\tilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\tilde{C}\geq t$ such that $\{a\}\times\tilde{C}\subseteq\tilde{Y}$. Hence $s\in S_{{\mathscr{F}}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)). Now, the conclusion of the lemma holds as $s$ runs over $E_{j}(t)$. ∎
@@
(3.8) $S_{{\mathscr{F}}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
So $S_{{\mathscr{F}}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is. The “In particular” part is easy to check once we have established the Zariski closedness of $S_{{\mathscr{F}}}(t)$. ∎

### v2

@@ 3. The Betti rank and Zariski closedness of the Betti strata
$\nu_{\mathrm{Betti},s}\colon T_{s}S\rightarrow T_{\nu(s)}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}},s})$
at each $s\in S({\mathbb{C}})$, and the Betti rank (1.5)
@@
The following formula gives the Betti rank:
@@
For each $t\geq 0$, the set $S^{\mathrm{Betti}}(t):=\{s\in S({\mathbb{C}}):\dim_{\nu(s)}(\nu(S)\cap{\mathscr{F}}_{\mathrm{Betti}})\geq t\}$ is Zariski closed in $S$. In particular, $r(\nu)=\dim S-\min\{t\geq 0:S^{\mathrm{Betti}}(t)=S\}$.
The proofs of Theorem 3.1 and Theorem 3.2 are simultaneous and follow the guideline of the first-named author’s [Gao20a] on the case when ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})$ is polarizable. A key ingredient for our proof is the mixed Ax–Schanuel theorem, which is used multiple times. In [Gao20a], the version for universal abelian varieties [Gao20b, Thm. 1.1] was used. In the current paper, we need the version for admissible VMHS independently proved by Chiu [Chi24] and Gao–Klingler [GK24]. We also invoke [BBT23] on the algebraicity of $\varphi(S)$, whose proof builds up on o-minimal GAGA, eases the notation for the proof.
@@ 3.1. Replacing $S$ by $\varphi(S)$
(3.3) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1}S_{\geq r}\cap\varphi^{-1}\left(S^{\prime}_{{\mathscr{F}}}(t-r)\right).$
@@ 3.2. Bi-algebraic system and Ax–Schanuel
    Let $\widetilde{Y}\subseteq{\mathscr{D}}$ be a complex analytic irreducible subset. The weakly special closure of $\widetilde{Y}$, denoted by $\widetilde{Y}^{\mathrm{ws}}$, is the smallest weak Mumford–Tate domain in ${\mathscr{D}}$ which contains $\widetilde{Y}$.
@@
The following Ax–Schanuel theorem for VMHS was independently proved by Chiu [Chi24] and Gao–Klingler [GK24]. We refer to Definition B.2 for the algebraic structure on ${\mathscr{D}}$.
@@
Let $\widetilde{Z}\subseteq u^{-1}(S)$ be a complex analytic irreducible subset. Then
(3.4) $\dim\widetilde{Z}^{\mathrm{Zar}}+\dim u(\widetilde{Z})^{\mathrm{Zar}}\geq\dim\widetilde{Z}^{\mathrm{ws}}+\dim\widetilde{Z},$
where $\widetilde{Z}^{\mathrm{ws}}$ is the smallest weak Mumford–Tate domain which contains $\widetilde{Z}$.
@@
Let $Y:=u(\widetilde{Z})^{\mathrm{Zar}}$. Let $\mathscr{Z}:=\{(z,y)\in\widetilde{Z}\times Y({\mathbb{C}}):u(z)=y\}$, then $\mathscr{Z}$ is a complex analytic irreducible subset of ${\mathscr{D}}\times_{\Gamma\backslash{\mathscr{D}}}Y^{\prime}$. The Zariski closure of $\mathscr{Z}$ in ${\mathscr{D}}\times Y$ is contained in $\widetilde{Z}^{\mathrm{Zar}}\times Y$, and $\dim\mathscr{Z}=\dim\widetilde{Z}$. Then (3.4) is a direct consequence of the mixed Ax–Schanuel theorem [GK24, Thm. 1.1] applied to $\mathscr{Z}$. ∎
@@
An irreducible algebraic subvariety $Y$ of $S$ is called weakly optimal if the following holds: $Y\subsetneq Y^{\prime}\subseteq S\Rightarrow\delta_{\mathrm{ws}}(Y)<\delta_{\mathrm{ws}}(Y^{\prime})$, for any $Y^{\prime}\subseteq S$ irreducible.
@@ 3.3. Applications of Ax–Schanuel
For each $t\geq 0$, $S^{\mathrm{Betti}}(t)$ is contained in the union of weakly optimal subvarieties $Y\subseteq S$ satisfying
@@
    $S^{\mathrm{Betti}}(t)$ is covered by the union of irreducible subvarieties $Y\subseteq S$ satisfying (3.6) (without requiring $Y$ to be weakly optimal);
@@
Let us prove (i). By Lemma 2.1, more precisely the first equality of (2.1), $S^{\mathrm{Betti}}(t)$ is covered by irreducible subvarieties $Y\subseteq S$ such that
$Y:=\overline{u(\{a\}\times\widetilde{C})}^{\mathrm{Zar}},$
for some complex analytic irreducible $\widetilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\widetilde{C}=t$ and some $a\in V({\mathbb{R}})$.
Apply mixed Ax–Schanuel in this context (Theorem 3.5 to $\{a\}\times\widetilde{C}$). Then we get
$\dim\overline{\{a\}\times\widetilde{C}}^{\mathrm{Zar}}+\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}+t.$
By Lemma B.4, $\overline{\{a\}\times\widetilde{C}}^{\mathrm{Zar}}=\{a\}\times\overline{\widetilde{C}}^{\mathrm{Zar}}$. Hence
$\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\overline{\widetilde{C}}^{\mathrm{Zar}}+t\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\widetilde{C}^{\mathrm{ws}}+t.$
The last inequality holds because $\overline{\widetilde{C}}^{\mathrm{Zar}}\subseteq\widetilde{C}^{\mathrm{ws}}$. So
$\dim Y\geq\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}-\dim\widetilde{C}^{\mathrm{ws}}+t.$
Now, to prove (i), it suffices to prove $\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$ and $\dim\widetilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$.
Let us prove $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$; the upshot is $\dim(\{a\}\times\widetilde{C})^{\mathrm{ws}}=\dim Y^{\mathrm{ws}}$. By definition of $Y$, we have $u(\{a\}\times\widetilde{C})\subseteq Y$. Hence $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})\subseteq Y^{\mathrm{ws}}$. On the other hand, $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$ is closed algebraic by Proposition 3.3, so $Y\subseteq u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$. So $Y^{\mathrm{ws}}\subseteq u((\{a\}\times\widetilde{C})^{\mathrm{ws}})$. Now we have established $u((\{a\}\times\widetilde{C})^{\mathrm{ws}})=Y^{\mathrm{ws}}$.
Similarly, we have $\dim\widetilde{C}^{\mathrm{ws}}=\dim[p](Y^{\mathrm{ws}})$. Hence, we are done for (i).
@@
There exist finitely many pairs $({\mathscr{D}}^{\prime}_{1},N_{1}),\ldots,({\mathscr{D}}^{\prime}_{k},N_{k})$, with each ${\mathscr{D}}^{\prime}_{j}$ a Mumford–Tate domain contained in ${\mathscr{D}}$ and $N_{j}$ a normal subgroup of $\mathrm{MT}({\mathscr{D}}^{\prime}_{j})$, such that the following holds. For each weakly optimal subvariety $Y\subseteq S$, $Y^{\mathrm{ws}}$ is the image of an $N_{j}({\mathbb{R}})^{+}$-orbit contained in ${\mathscr{D}}^{\prime}_{j}$ under $u\colon{\mathscr{D}}\rightarrow\Gamma\backslash{\mathscr{D}}$ for some $j\in\{1,\ldots,k\}$.
The same statement (with $Y^{\mathrm{ws}}$ replaced by $\varphi(Y)^{\mathrm{ws}}$) still holds without replacing $S$ by $\varphi(S)$.
@@
$S^{\mathrm{Betti}}(t)\subseteq\bigcup\nolimits_{j=1}^{k}E_{j}(t).$
@@
Let $t\geq 0$. By Proposition 3.7, $S^{\mathrm{Betti}}(t)$ is covered by weakly optimal $Y\subseteq S$ such that $\dim Y\geq\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})+t$. Then by Proposition 3.8, $Y^{\mathrm{ws}}$ is a fiber $[p_{N_{j}}]$ for some $j\in\{1,\ldots,k\}$, and hence $\dim Y^{\mathrm{ws}}-\dim[p](Y^{\mathrm{ws}})=\frac{1}{2}\dim(V\cap N_{j})$. So $\dim Y\geq\frac{1}{2}\dim(V\cap N_{j})+t$. So $Y\subseteq E_{j}(t)$ because $[p_{N_{j}}](Y)$ is a point. ∎
@@
Let us prove $\leq$”. By (2.6) we have $r(\nu)=\max_{\widetilde{s}\in\widetilde{S}}(\mathrm{d}\widetilde{\varphi}_{V})_{\widetilde{s}}$. Hence $r(\nu)\leq\dim\varphi_{/N}(S)+\frac{1}{2}\dim(V\cap N)$ for any normal subgroup $N\lhd\mathbf{G}$.
Let us prove $\geq$”. Let $t=\dim S-r(\nu)$. Then $S^{\mathrm{Betti}}(t)$ contains a non-empty open subset of $S^{\mathrm{an}}$. By (3.3) and Proposition 3.9 (which should be applied to $S^{\prime}_{{\mathscr{F}}}(t-r)$ for each $10\leq r\leq t-$ 1), we have
$S^{\mathrm{Betti}}(t)\subseteq S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
@@
If $S=S_{\geq t}$, then $\geq$” holds already for $N=\{1\}$.
@@ 3.5. Zariski closedness of the degeneracy loci
For each $t\geq 0$ and each $j\in\{1,\ldots,k\}$, we have $E_{j}(t)\subseteq S^{\mathrm{Betti}}(t)$.
@@
For each $s\in E_{j}(t)$, by definition there exist an irreducible $\widetilde{Y}\subseteq u^{-1}(S)\cap{\mathscr{D}}^{\prime}_{j}$ such that $s\in u(\widetilde{Y})$, $\dim\widetilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, and that $\widetilde{Y}$ is contained in a fiber of the quotient ${\mathscr{D}}^{\prime}_{j}\rightarrow{\mathscr{D}}^{\prime}_{j}/N_{j}$. The last condition implies that $\widetilde{Y}\subseteq N_{j}({\mathbb{R}})^{+}x$ for some $x\in{\mathscr{D}}^{\prime}_{j}$. Hence by the discussion above, $\widetilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$ for a fixed $v\in V({\mathbb{R}})$. Now that $\dim_{{\mathbb{C}}}\widetilde{Y}\geq\frac{1}{2}\dim(V\cap N_{j})+t$, the following property holds: For each $(a,x_{0})\in\widetilde{Y}\subseteq((V\cap N_{j})({\mathbb{R}})+v)\times{\mathscr{D}}_{0}$, there exists a complex analytic subset $\widetilde{C}\subseteq{\mathscr{D}}_{0}$ with $\dim\widetilde{C}\geq t$ such that $\{a\}\times\widetilde{C}\subseteq\widetilde{Y}$. Hence $s\in S^{\mathrm{Betti}}(t)$ by Lemma 2.1 (more precisely, the first equality in (2.1)). Now, the conclusion of the lemma holds as $s$ runs over $E_{j}(t)$. ∎
@@
(3.8) $S^{\mathrm{Betti}}(t)=S_{\geq t}\cup\bigcup\nolimits_{0\leq r\leq t-1,~{}1\leq j\leq k}S_{\geq r}\cap\varphi^{-1}\left(E_{j}(t-r)\right).$
So $S^{\mathrm{Betti}}(t)$ is Zariski closed in $S$ because each member in the union on the right-hand side is. The “In particular” part is easy to check once we have established the Zariski closedness of $S^{\mathrm{Betti}}(t)$. ∎